OpenStudy (anonymous):

Solve -4√x+2 + 3 = -9 a. x=-1 b. x= 7 c. x= 31 d. x= 46 I thought 7... but that doesn't seem to work

OpenStudy (anikhalder):

take \[\sqrt{x}=p\] So, equation stands as, -4p+2+3=-9 4p=14 p=3.5 Hence \[\sqrt{x}=3.5\] Squaring we get x=3.5*3.5 So, x=12.25 which is not among the answers Will there be an x after 2?

OpenStudy (anonymous):

no, i dont think so.. i don't get where you got the 14 from though.

OpenStudy (anonymous):

nevermind, it's the 2+3+9 .

OpenStudy (anonymous):

See i had gotten the 12.25 as well, but seeing as that is not one of the answers, i dont get it.

OpenStudy (anikhalder):

hmm..this is pretty confusing..:)

OpenStudy (anonymous):

hmm, i can't figure it out /:

OpenStudy (anikhalder):

Lets plug in each of the values a. definitely not it as it gives a complex root b. plugging x=7 gives a value of -5.5 instead of 9. Cross it out too. c. LHS will give a way negative value instead of 9. Wrong too d. Same fate as c. Cross it out too. We should recheck the question itself shouldn't we?

OpenStudy (anonymous):

I mean, that's all it gives me. exactly that. So , is there any other way of solving this?

OpenStudy (anikhalder):

Hmm..No idea...Am sorry..cannot help..will try and figure out something else..nothing coming to my mind now..

OpenStudy (zehanz):

I think the question is different: \(-4\sqrt{x+2}+3=-9\). Then \(-4\sqrt{x+2}=-12\), so \(\sqrt{x+2}=3\). Not there are only two steps left to get your answer!

OpenStudy (zehanz):

@brazukinhaa: seems like your original thought was not so bad...

OpenStudy (anikhalder):

@ZeHanz ....You are right..:)

OpenStudy (zehanz):

Typing -4√(x+2) + 3 = -9 would have made it much clearer. It is because in a book or on a webpage these numbers are nicely typesetted like \(-4\sqrt{x+2}+3=-9\), but people tend to forget to type brackets if they don't use the equation editor or LATEX commands.

OpenStudy (anonymous):

Sorry i had gone out. Okay, so i'm still not understanding which of the numbers would be the answer. but i see what you're saying. I couldn't write it like that however, i couldn't find a radical that would remain over the x+2 @ZeHanz

OpenStudy (zehanz):

OK, we've come to: \(-4\sqrt{x+2}+3=-9\), so \(-4\sqrt{x+2}=-12\) (subtracted 3 from both sides) \(\sqrt{x+2}=3\) (divided both sides by -4) Now, don't pay attention to the number x+2 yet. Just ask yourself: the square root of what number is 3? Well, 9 of course! So the number under the radical must be 9: \(x+2=9\) and this means that \(x=7\).

OpenStudy (anonymous):

Yeah, that's what i had originally though, because 7+2 = 9 and then the square root being 3. I think i get it now. Thanks so muchh! (:

OpenStudy (zehanz):

YW!